# arXiv:2407.08738v2[cond-mat.quant-gas] 28 Feb 2026

### v2

@@ Abstract
We report the numerical observation of a far-from-equilibrium equation of state (EOS) in the Gross-Pitaevskii model. We first show that the momentum distribution of the turbulent cascade is well described by wave-turbulent kinetic theory in the appropriate limits. Calculating the energy and particle fluxes $\Pi_{\varepsilon}(k)$ and $\Pi_{N}(k)$, we show that the turbulent state possesses the hallmarks of a direct energy cascade. Building on this, we show that the GP model encodes a universal EOS in the form of a relationship between the turbulent cascade’s momentum distribution amplitude $n_{0}$ and the energy flux $\epsilon$ in the steady state. We find that in our regime of ‘mixed’ turbulence - where both vortices and waves play a significant role - $n_{0}\propto\epsilon^{0.67(2)}$, a result that is not captured by any existing theory of turbulence but that agrees with a recent experimental measurement for large energy fluxes. Finally, we find that the concept of quasi-static thermodynamic processes between equilibrium states extends to far-from-equilibrium steady states.
@@
The system is initialized in the ground state of Eq. (1) including $V_{\text{box}}(\bm{r})$, which corresponds to a Bose-Einstein condensate with a nearly uniform density, except near the boundary of the box. For relevance, we use numbers typical of recent experiments [14]: a box radius $R=15\,\upmu{\rm m}$ and length $L=50\,\upmu{\rm m}$, an initial atom number $N(t=0)=2\times 10^{5}$, and $a/a_{0}$ between $25$ and $400$ (where $a_{0}$ is the Bohr radius). The natural energy scale of the system $\zeta\equiv gn(t=0)$ therefore varies between $k_{\textrm{B}}\times 1.2\,{\rm nK}$ and $k_{\textrm{B}}\times 19\,{\rm nK}$, the corresponding timescale $\tau\equiv h/\zeta$, between $2.6\,\textrm{ms}$ and $42\,\textrm{ms}$, and the natural lengthscale $\xi\equiv\sqrt{2m\zeta}/\hbar$, between $0.6\,\upmu{\rm m}$ and $2.3\,\upmu{\rm m}$ (using the mass $m$ of 39K); $n=N/V$ is the average density, where $V=\pi R^{2}L$ is the box volume. The gas is driven at a frequency $\omega/(2\pi)$ that matches the resonance of the lowest-lying Bogoliubov excitation [39]; for our set of parameters, $\omega/(2\pi)$ ranges from $5$ to $20\,\textrm{Hz}$.
@@
We systematically studied the momentum-resolved cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}\ln{[k]}$ across interaction strengths (different $\xi$) and find that $\gamma(k\xi)$ is in excellent agreement with the $4$-wave WWT prediction $3+1/(3\ln^{1/3}[k/k_{\textrm{F}}])$, provided that $k_{\textrm{F}}$ is replaced by $k_{0}=1.64(2)k_{\xi}$ (see End Matter and Fig. 5). In other words, the isotropic $4$-wave cascade’s effective injection scale is set by the GPE scale $1/\xi$ instead of the physical injeciton scale $k_{\textrm{F}}$. However, as was already presumed experimentally [15, 14], we find that a power law $n(k)\propto k^{-\gamma_{0}}$ with an effective exponent $\gamma_{0}=3.5$ accurately captures all the relevant details, since the variation of the logarithmic correction for $\gamma(k)$ is small over the relevant momentum range  [^6] Note that [20] speculated that the steeper effective exponent could be due to the logarithmic correction. However, this work assumes that there is no BEC, but the presence of the BEC is crucial for understanding the effective injection scale $k_{0}\propto k_{\xi}$, hence [20] cannot make quantitative comparisons to the experiment [15]. (see also a discussion in  [36]). From now on, we fix $n(k)$ to this power law and define the amplitude of the power law $n_{0}\equiv N_{k}k^{3}(k\xi)^{\gamma_{0}-3}$.
@@
Energetics of the turbulent cascade. We next turn to the calculation of the energy-density input and dissipation rates. As energy is injected into the system only by the forcing $V_{\text{drive}}$, the energy input rate can be calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ where $F\equiv-\hat{z}\cdot\nabla V_{\text{drive}}$, $\hat{z}$ is the unit vector along $z$, $v$ is the center-of-mass velocity of the gas and the averaging $\langle\cdot\rangle$ is performed over a drive period. The energy is dissipated at high momenta solely by particles with momenta $k>k_{\textrm{D}}$ leaving the trap. Hence, in previous experimental works [16, 14], the energy dissipation rate was reasonably assumed to be $\dot{N}U_{\textrm{D}}$ where $\dot{N}$ is the particle loss rate. However, as shown in Fig. 2(a) for three different $U_{\textrm{s}}$, the energy-density injection rate $\epsilon_{\textrm{in}}$ is consistently higher than $\dot{N}U_{\textrm{D}}/V$. The reason for $\epsilon_{\textrm{in}}>\dot{N}U_{\textrm{D}}/V$ stems from the shape of the dissipation spectrum $D_{\varepsilon}(k)$ (see [36] for a formal definition): calculating the energy-density dissipation rate as $\dot{N}U_{\textrm{D}}/V$ assumes that the dissipation happens exactly at energy $U_{\textrm{D}}$. As shown in the inset of Fig. 2(a) for $U_{\textrm{s}}=\zeta$, $D_{\varepsilon}(k)$ has a sharp onset at $k_{\textrm{D}}$, but has a significant tail above $k_{\textrm{D}}$. The average dissipation momentum is $\langle k_{\textrm{diss}}\rangle\equiv\int kD_{\varepsilon}(k)\textrm{d}k/\int D_{\varepsilon}(k)\textrm{d}k\approx 1.15k_{\textrm{D}}$  [^7] A simple geometric picture is that particles hitting the box at non-normal incidence can remain trapped for some time even if their kinetic energy $\hbar^{2}k^{2}/(2m)>U_{\textrm{D}}$., showing that the energy dissipated from the system per unit time is $\approx 1.32\dot{N}U_{\textrm{D}}$. This is in excellent agreement with $\alpha\equiv\epsilon_{\textrm{in}}/(\dot{N}U_{\textrm{D}})$  [^8] Even though $\alpha\neq 1$ means that the energy-density flux is not $\dot{N}U_{\textrm{D}}/V$, the fact that $\alpha$ has only a very weak dependence on $a$ ($\propto a^{0.08(1)}$) and is independent of $U_{\textrm{s}}$ suggests that the determination of the energy flux via $\dot{N}U_{\textrm{D}}/V$, as done in [16, 14], will only introduce a small, presumably system-specific, systematic error. [see the End Matter for a systematic study of $\alpha$ versus interaction strength and $U_{\textrm{s}}$, showing that $\alpha\approx 1.3$].
@@
Having verified that the energy input rate (at low $k$) and dissipation rate (at high $k$) are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [36] for the formal definition of the fluxes). As shown in Fig. 2(c) in the steady state, $\Pi_{\varepsilon}$ is momentum independent in the inertial range, a certain range of momenta where neither forcing nor dissipation takes place; this flux transports the energy injected at low $k$ to higher $k$, up to $k_{\textrm{D}}$. Importantly, $\Pi_{\varepsilon}$ in the inertial range is equal to $\epsilon_{\textrm{in}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^9] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the upper panel of Fig. 2(c)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(c)]  [^10] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [56].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
@@
Conclusions. We numerically investigated the properties of a turbulent cascade arising in the GP model when a box trapped gas is periodically driven, and showed that it can be described by a universal EOS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EOS is inconsistent with any existing theory of turbulence, posing a new theoretical challenge. Furthermore, the comparison of our classical-field simulations to experimental measurements provides valuable benchmarks for testing the validity of classical-field theories in far-from-equilibrium scenarios. Namely, we find that the GP model provides valuable insight into the experimental results of [14], even if it fails to fully capture the equation of state – indeed, the observed $na^{3}$ scaling is incompatible with the universal GP framework (see also [36]). This partial success is nevertheless striking given that there is no firm theoretical justification for its applicability in this turbulent regime. Understanding the foundations of this success, and developing refined approaches that incorporate quantum effects, remain important challenges for the field. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.

### v3

# An Equation of State for Turbulence in the Gross-Pitaevskii Model

@@ Abstract
We report the numerical observation of a far-from-equilibrium equation of state (EOS) in the Gross-Pitaevskii (GP model. We first show that the momentum distribution of the turbulent cascade is well described by wave-turbulent kinetic theory in the appropriate limits. Calculating the energy and particle fluxes $\Pi_{\varepsilon}(k)$ and $\Pi_{N}(k)$, we show that the turbulent state possesses the hallmarks of a direct energy cascade. Building on this, we show that the GP model encodes a universal EOS in the form of a relationship between the turbulent cascade’s momentum distribution amplitude $n_{0}$ and the energy flux $\epsilon$ in the steady state. We find that in our regime of ‘mixed’ turbulence - where both vortices and waves play a significant role - $n_{0}\propto\epsilon^{0.67(2)}$, a result that is not captured by any existing theory of turbulence but that agrees with a recent experimental measurement for large energy fluxes. Finally, we find that the concept of quasi-static thermodynamic processes between equilibrium states extends to far-from-equilibrium steady states.
@@
The system is initialized in the ground state of Eq. (1) including $V_{\text{box}}(\bm{r})$, which corresponds to a Bose-Einstein condensate with a nearly uniform density, except near the boundary of the box. For relevance, we use numbers typical of recent experiments [14]: a box radius $R=15\,\upmu{\rm m}$ and length $L=50\,\upmu{\rm m}$, an initial atom number $N(t=0)=2\times 10^{5}$, and $a/a_{0}$ between $25$ and $400$ (where $a_{0}$ is the Bohr radius). The natural energy scale of the system $\zeta\equiv gn(t=0)$ therefore varies between $k_{\textrm{B}}\times 1.2\,{\rm nK}$ and $k_{\textrm{B}}\times 19\,{\rm nK}$, the corresponding timescale $\tau\equiv h/\zeta$, between $2.6\,\textrm{ms}$ and $42\,\textrm{ms}$, and the natural lengthscale $\xi\equiv\hbar/\sqrt{2m\zeta}$, between $0.6\,\upmu{\rm m}$ and $2.3\,\upmu{\rm m}$ (using the mass $m$ of 39K); $n=N/V$ is the average density, where $V=\pi R^{2}L$ is the box volume. The gas is driven at a frequency $\omega/(2\pi)$ that matches the resonance of the lowest-lying Bogoliubov excitation [39]; for our set of parameters, $\omega/(2\pi)$ ranges from $5$ to $20\,\textrm{Hz}$.
@@
We systematically studied the momentum-resolved cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}\ln{[k]}$ across interaction strengths (different $\xi$) and find that $\gamma(k\xi)$ is in excellent agreement with the $4$-wave WWT prediction $3+1/(3\ln[k/k_{\textrm{F}}])$, provided that $k_{\textrm{F}}$ is replaced by $k_{0}=1.64(2)k_{\xi}$ (see End Matter and Fig. 5). In other words, the isotropic $4$-wave cascade’s effective injection scale is set by the GPE scale $1/\xi$ instead of the physical injeciton scale $k_{\textrm{F}}$. However, as was already presumed experimentally [15, 14], we find that a power law $n(k)\propto k^{-\gamma_{0}}$ with an effective exponent $\gamma_{0}=3.5$ accurately captures all the relevant details, since the variation of the logarithmic correction for $\gamma(k)$ is small over the relevant momentum range  [^6] Note that [20] speculated that the steeper effective exponent could be due to the logarithmic correction. However, this work assumes that there is no BEC, but the presence of the BEC is crucial for understanding the effective injection scale $k_{0}\propto k_{\xi}$, hence [20] cannot make quantitative comparisons to the experiment [15]. (see also a discussion in  [36]). From now on, we fix $n(k)$ to this power law and define the amplitude of the power law $n_{0}\equiv N_{k}k^{3}(k\xi)^{\gamma_{0}-3}$.
@@
Energetics of the turbulent cascade. We next turn to the calculation of the energy-density input and dissipation rates. As energy is injected into the system only by the forcing $V_{\text{drive}}$, the energy input rate can be calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ where $F\equiv-\hat{z}\cdot\nabla V_{\text{drive}}$, $\hat{z}$ is the unit vector along $z$, $v$ is the center-of-mass velocity of the gas and the averaging $\langle\cdot\rangle$ is performed over a drive period. The energy is dissipated at high momenta solely by particles with momenta $k>k_{\textrm{D}}$ leaving the trap. Hence, in previous experimental works [16, 14], the energy dissipation rate was reasonably assumed to be $\dot{N}U_{\textrm{D}}$ where $\dot{N}$ is the particle loss rate. However, as shown in Fig. 2(a) for three different $U_{\textrm{s}}$, the energy-density injection rate $\epsilon_{\textrm{in}}$ is consistently higher than $\dot{N}U_{\textrm{D}}/V$. The reason for $\epsilon_{\textrm{in}}>\dot{N}U_{\textrm{D}}/V$ stems from the shape of the dissipation spectrum $D_{\varepsilon}(k)$ (see [36] for a formal definition): calculating the energy-density dissipation rate as $\dot{N}U_{\textrm{D}}/V$ assumes that the dissipation happens exactly at energy $U_{\textrm{D}}$. As shown in the inset of Fig. 2(a) for $U_{\textrm{s}}=\zeta$, $D_{\varepsilon}(k)$ has a sharp onset at $k_{\textrm{D}}$, but has a significant tail above $k_{\textrm{D}}$. The average dissipation momentum is $\langle k_{\textrm{diss}}\rangle\equiv\int kD_{\varepsilon}(k)\textrm{d}k/\int D_{\varepsilon}(k)\textrm{d}k\approx 1.15k_{\textrm{D}}$  [^7] A simple geometric picture is that particles hitting the box at non-normal incidence can remain trapped for some time even if their kinetic energy $\hbar^{2}k^{2}/(2m)>U_{\textrm{D}}$., showing that the energy dissipated from the system per unit time is $\approx 1.32\dot{N}U_{\textrm{D}}$. This is in excellent agreement with $\alpha\equiv\epsilon_{\textrm{in}}/(\dot{N}VU_{\textrm{D}})$  [^8] Even though $\alpha\neq 1$ means that the energy-density flux is not $\dot{N}U_{\textrm{D}}/V$, the fact that $\alpha$ has only a very weak dependence on $a$ ($\propto a^{0.08(1)}$) and is independent of $U_{\textrm{s}}$ suggests that the determination of the energy flux via $\dot{N}U_{\textrm{D}}/V$, as done in [16, 14], will only introduce a small, presumably system-specific, systematic error. [see the End Matter for a systematic study of $\alpha$ versus interaction strength and $U_{\textrm{s}}$, showing that $\alpha\approx 1.3$].
@@
Having verified that the energy input rate (at low $k$) and dissipation rate (at high $k$) are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [36] for the formal definition of the fluxes). As shown in Fig. 2(b) in the steady state, $\Pi_{\varepsilon}$ is momentum independent in the inertial range, a certain range of momenta where neither forcing nor dissipation takes place; this flux transports the energy injected at low $k$ to higher $k$, up to $k_{\textrm{D}}$. Importantly, $\Pi_{\varepsilon}$ in the inertial range is equal to $\epsilon_{\textrm{in}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^9] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the left panel of Fig. 2(b)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(b)]  [^10] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [56].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
@@
Conclusions. We numerically investigated the properties of a turbulent cascade arising in the GP model when a box-trapped gas is periodically driven, and showed that it can be described by a universal EOS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EOS is inconsistent with any existing theory of turbulence, posing a new theoretical challenge. Furthermore, the comparison of our classical-field simulations to experimental measurements provides valuable benchmarks for testing the validity of classical-field theories in far-from-equilibrium scenarios. Namely, we find that the GP model provides valuable insight into the experimental results of [14], even if it fails to fully capture the equation of state – indeed, the observed $na^{3}$ scaling is incompatible with the universal GP framework (see also [36]). This partial success is nevertheless striking given that there is no firm theoretical justification for its applicability in this turbulent regime. Understanding the foundations of this success, and developing refined approaches that incorporate quantum effects, remain important challenges for the field. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.